# Tackling Cyberattacks through AI-based Reactive Systems: A Holistic Review and Future Vision

### v1

@@ III-B Algorithm
  • Neural Networks (NN): Self-taught learning algorithm inspired by the functioning of the biological nervous system, particularly neurons in the human brain [72]. It consists of interconnected nodes that interact and process information to be able to recognize patterns, classification, and regression, among others. Figure 4 displays an example of NN, in which three layers are detailed. The first layer is in charge of receiving the data, in the hidden layer particular processes are carried out, which can finally be seen in the output layer that delivers the desired result. This type of network can be used in reactive systems to adapt to changing threats by the countermeasure selection system. Furthermore, it could be used for threat classification to select the most optimal countermeasures to be applied.
    Figure 4: Graphical Neural Network example
  • A* algorithm: A search algorithm used in artificial intelligence to find the shortest path or optimal route between an initial node and a target node in a network. This algorithm uses a combination of heuristic search and uniform cost search to determine the most efficient path [73]. In a reactive system it could be used to find optimal paths in order to select countermeasures based on defined costs. Figure 5 shows a graphical example of the search for an optimal path to solve a problem. In a threat response framework, this algorithm could be used to analyze network or countermeasures graphs, enhancing the system capability to calculate an optimal path.
    Figure 5: Graphical A* algorithm example
  • Genetic Algorithms: Optimization and search techniques based on the theory of biological evolution [74]. These algorithms are used to find approximate solutions to optimization problems. The process resembles the evolutionary process, where candidate solutions are combined, mutated, and selected to form new solutions. Using concepts such as natural selection, crossover, and mutation to improve the solutions. In threat response systems can be used, for example, by evolving a population of potential feature countermeasures in helping to identify the optimal set of security measures.
@@ III-C Strategy Objectives
  • Minimizing risk: This objective reflects the intention to reduce exposure to potential risks or adverse events. Such an objective indicates a significant concern to manage any uncertainties that may exist and to ensure the stability and reliability of the architecture.
  • Threat detection: This objective represents the need to identify potential threats. It emphasises the importance of proactive measures and early detection of possible vulnerabilities. It is the objective presented in the IDS.
  • Minimising risk and cost: The combination of these objectives means a dual focus on cost-effectiveness and risk mitigation. It implies a strategy aimed at balancing economic considerations with security concerns.
  • Countermeasure standard proposal: This objective indicates the intention to develop and propose standardised countermeasures. At the same time, it suggests a proactive role in the configuration of best practices and defensive strategies in the field.
@@ III-D Measurements and results
  • Positive rate: Often referred to as a true positive rate. Is a performance metric that measures the proportion of true positive predictions out of all actual positive instances  [76]. Equation 1 shows the mathematical expression of this metric. Where true positives represent instances that a model correctly recognizes or predicts as belonging to the attribute being tested. And, false negatives represent instances of the positive class that were incorrectly identified as negative.
    $True\_positive\_rate=\frac{True\_positives}{True\_positives+False\_Negatives}$ (1)
@@ III-E Use Case
  • Critical Infrastructure: This value is applicable in the field of critical infrastructures. Critical infrastructures refer to the fundamental systems, assets, facilities, and networks that are vital for the functioning of a society and its economy [77, 78]. Such a value implies the relevance and effectiveness of the proposed system in safeguarding and optimising the essential components of the infrastructure.
  • Internet of Things (IoT): This value emphasises the adaptability and functionality of a proposed system into the interconnected IoT ecosystem. IoT refers to a vast network of interconnected devices, objects, and systems that can communicate and exchange data with each other over the internet [79, 80].
@@ III-F Features Summary
  • Strategy Objectives where seven values are recognized: Minimizing risk, Vulnerability assessment, Minimising risk and cost, Threat Detection, Calculation of cost-damage probabilities in attacks, Mitigate attack and Countermeasure standard proposal.
@@ IV Survey
For each of the papers a description of the research proposal is conducted. Moreover, an analysis is provided highlighting the advantages and/or disadvantages of each of the approaches. Furthermore, this analysis is based on the features selected in Section III, ensuring that each paper will be discussed under the same conditions in terms of the proposed features.
@@ IV-A Cybersecurity risk analysis model using fault tree analysis and fuzzy decision theory [81]
In [81], the authors present a model that integrates Fault Tree Analysis (FTA) [82], decision theory [33], and fuzzy theory [83] to address two key objectives. FTA provides a structured and systematic way to identify and evaluate potential system failures by modeling the relationships between various components or events within a system. Decision theory is a field that encompasses various methods and tools for making rational and informed decisions. Fuzzy theory is a mathematical framework that allows for the representation of uncertainty and vagueness in a systematic manner. On the one hand, the model allows to identify the underlying causes of failures in cyber-attack prevention. On the other hand, it enables to assess the vulnerability of a given system.
Initially, the FTA methodology is employed to establish a failure event associated with a potential cyber-attack and systematically trace its underlying influential factors. Particularly, the primary objective of FTA is to identify the minimal cut set, which represents the minimum combination of basic events leading to the occurrence of the top event, for example the combination: “CyberAttack - Fault Information System -Unauthorized Access”. By analyzing these cut sets, appropriate actions can be prioritized and implemented to prevent the manifestation of the top event and identify vulnerabilities within the system. To comprehensively assess the complexity, magnitude, and impact of a cyber-attack, this research introduces the utilization of fuzzy theory and decision-making techniques under conditions of uncertainty.
@@
The third phase, i.e., the definition of criteria, focuses on addressing the uncertainties involved in risk analysis in the context of cybersecurity. The paper proposes the use of models presented in [84] and emphasizes the need to consider multiple criteria for evaluating the consequences of a cybersecurity solution. Specifically, the two criteria used in this paper are financial losses and time for restoration. In this sense, the importance of considering cost implications and the ability to repair or replace damaged services quickly is highlighted. The fourth phase focuses on the fuzzy assessment of potential accidents. During this stage, managerial experts evaluate alternative options using the identified criteria. Finally, the last phase involves aggregating all of the criteria used and ordering the alternatives, according to the magnitude of their consequences and using custom equations. These equations are designed for representing maximum risk using an estimated consequence of actions under nature states. The purpose of this representation is used to adopt countermeasures that minimise the risk identified.
@@
The initial focus of the research is on the data representation used for the countermeasure selection in the cybersecurity context. To achieve this objective, the authors define specific information objects that serve as input for the process. These objects are vulnerabilities, products, weaknesses, exploits, attack patterns, configurations, and countermeasures. The relationships between these objects are established through the analysis of open security databases such as Snyk [86] or Mend [87]. In this manner, events can be associated with these objects to enhance their representation and enable more accurate utilization of the information. The data collected by the system can be categorized into two types: input data and output data. Input data encompasses information obtained from external sources, in particular vulnerabilities, code or architecture weaknesses, and attack patterns. It also includes data generated through network analysis, such as network events, configurations, countermeasures, and security policies. In contrast, output data refers to the information generated as a result of attack modeling and analysis, including, AGs, security metrics, impact assessment, selected countermeasures, and identified weaknesses. The system utilizes both input and output data to enhance its understanding of the security landscape and enable effective decision-making.
Following the data representation, the research introduces the proposed model for analyzing attacks. This model adheres to the subsequent procedures: firstly, an AG is constructed by considering existing and potential vulnerabilities. Next, the actions executed by malicious actors are identified using logs and alerts. Such procedure leads to the generation of an attack subgraph that represents the sequence of possible actions performed by the malicious actor. Furthermore, the model includes the incorporation of potential attack scenarios and corresponding countermeasures, as well as the calculation of security metrics for risk analysis. The resulting information is presented as events and data, enabling the selection of appropriate countermeasures and responses.
@@ IV-C Heuristic Approach Towards Countermeasure Selection using Attack Graphs [89]
In [89], an approach of heuristic search is followed to select the optimal countermeasure deployment under a given budget constraint. The approach implements AGs to model the risk present in the system. In order to do this, the proposed system considers vulnerability exploitation consequences and examines potential attacker lateral movements. Additionally, it represents the network’s physical topology as well as the vulnerabilities of the network protocols to minimize the system risk within the budgetary limitations. The solution relies on the shortest path problem where the A* heuristic algorithm [73] is used to find the most optimal set of countermeasures efficiently. The novelty of the article relies upon using a particular risk equation, which indicates the presented risk of exploiting an asset provided by CVSS scores, together with the associated countermeasure, thus allowing the evaluation of the risk of the system without regenerating the AG, saving computational time. MulVAL [90] is used as a security analyze based on logic, serving as a tool for the analysis of AGs. This tool uses the Prolog [91] language to represent the pre and post-conditions for exploiting a vulnerability and then identify the necessary countermeasure. The authors define countermeasures as security products with two attributes, on the one hand, the defense mechanism they implement, such as firewall or Antivirus, and, on the other hand, the set of mitigation actions, for example disabling/enabling ports. The proposed system works as follows. As a starting point, a risk assessment is carried out using the AGs. Subsequently, relevant countermeasures are identified by matching them with mitigation actions at the nodes of the AG, and identifying where the actions would be executed in the network. As a next step, the risk calculation is performed, using the risk equations that come from the preprocessing phase of the AGs. Once the risk has been calculated, the A* algorithm is used to obtain the shortest path, which in this case corresponds to the one with the lowest risk, considering the cost, thus identifying an optimal countermeasure plan. Finally, an evaluation of the proposal is carried out. For this purpose, costs are assigned as an effort for the deployment of each countermeasure. The algorithm will use a general budget and will reduce the risk, reducing the budget as an expense after the deployment of the corresponding countermeasure. It can be seen that the risk of the system decreases as the budget increases, but also the execution time of the algorithm increases, because additional countermeasures must be considered, which might incur higher costs.
@@
As mentioned before, this paper presents a methodology based on AGs for the purpose of risk minimization of a system. For this purpose, an A* algorithm is used for the selection of countermeasures. Analyzing the paper, a disadvantage identified concerns the limitations that conflict with the policies of the organization in which the system is implemented. In other words, the algorithm does not take into account certain limitations when selecting countermeasures. That is, in the risk graph shown in the paper, the higher the budget, the more the risk is reduced, even to 0%. This is however not useful in the real world, since the algorithm could propose countermeasures that disable relevant services or communications. Nevertheless, the use of MulVAL for AGs security analysis is an interesting point. Furthermore, it alludes to establishing a maximum cost, on which the heuristic-based algorithm is built, which increases the realism of the design. Therefore, it is easy to spot that this is a very interesting proposal that, once the possibility of taking into account certain requirements is added, can be very relevant.
@@ IV-D Introducing Deep Learning Self-Adaptive Misuse Network Intrusion Detection Systems [92]
Due to the large number of interconnected devices that may exist in a network, the ability to handle their capabilities is tipically beyond human limits. In the same way, IDSs are unable to automatically adapt to the many changes that occur in their environment, which reduces their efficiency. For example, by inserting a new device into the network it may generate services or communications or even vulnerabilities never seen in the network before. To solve the aforementioned challenges, in [92], Deep Learning (DL) is proposed to identify unknown threats in unknown environments. In particular, self-taught learning methodology is used, a ML methodology that can train itself using unlabeled data to improve a supervised classification problem. This methodology is integrated within the MAPE-K structure [93], which is based on the steps Monitor (M), Analyze (A), Plan (P), and Execute(E) while using a Knowledge (K) base as support. The proposed system is based on environmental states and network flows. Authors claim that network changes can be seen as potential incidents that may contain attacks or vulnerabilities. For this purpose, network flows are used, which, by possesing relevant information, can identify unknown and unseen network flows that may indicate the presence of previously unknown attacks.
In the monitor phase, the sensors of the system are coordinated to acquire the basic knowledge of the network, as well as network mappers [94] generate a network inventory. The idea is to identify the network characteristics alteration events that require activating the IDS. In parallel, network sniffers are used to capture network flows through the detection engine of the IDS. In this phase, network traffic is stored in the Knowledge base for later phases. In the Analyze phase, an analyzer leverages network audit tools to generate structured network flows from network traffic that can bring various features. These features are fed into the IDS engine, which will in turn use a supervised model for detection purposes and, on the other hand, it will be used as unlabeled data for the self-taught learning adaptive process.
Up to this point, the Monitor and the Analyser identified environmental changes in the network, while the Knowledge component consolidated the network flows generated at the time the changes occurred. At this moment, the IDS may face unknown network traffic, which is why the planning phase undertakes to utilize unsupervised feature learning techniques. This technique is a NN that applies backpropagation called sparse autoencoder [95] to learn sparse representations of the input data. The output of the planning phase is a feedforward autoencoder that is used to reconstruct the initial set of labeled data into a new training dataset for a supervised learning algorithm. After training the new model, the old model, which due to environmental changes had started to have efficiency problems, can be replaced.

### v2

@@ III-B Algorithm
  • Neural Networks (NN): Self-taught learning algorithm inspired by the functioning of the biological nervous system, particularly neurons in the human brain [74]. It consists of interconnected nodes that interact and process information to be able to recognize patterns, classification, and regression, among others. Figure 6 displays an example of NN, in which three layers are detailed. The first layer is in charge of receiving the data, in the hidden layer particular processes are carried out, which can finally be seen in the output layer that delivers the desired result. This type of network is useful in reactive systems, which adapt to evolving threats. By assimilating patterns from large sets of trace data, neural networks excel in anomaly detection, allowing the identification of suspicious activities. Furthermore, they could be used in threat classification to select the most optimal countermeasures to be applied.
    Figure 6: Graphical representation of the architecture of a standard Neural Network.
  • A* algorithm: A search algorithm used in AI to find the shortest path or optimal route between an initial node and a target node in a network. This algorithm uses a combination of heuristic search and uniform cost search to determine the most efficient path [75]. In a reactive system it could be used to find optimal paths in order to select countermeasures based on defined costs. Offering a powerful path finding, efficiently navigating through complex networks to identify the most effective countermeasure selection strategy. Figure 7 shows a graphical example of the search for an optimal path to solve a problem. In a threat response framework, this algorithm could be used to analyze network or countermeasures graphs, enhancing the system capability to calculate an optimal path.
    Figure 7: Graphical A* algorithm example of a path finding solution.
  • Genetic Algorithms: Optimization and search techniques based on the theory of biological evolution [76]. These algorithms are used to find approximate solutions to optimization problems. The process resembles the evolutionary process, where candidate solutions are combined, mutated, and selected to form new solutions. Using concepts such as natural selection, crossover, and mutation to improve the solutions. In threat response systems can be used, for example, by evolving a population of potential feature countermeasures in helping to identify the optimal set of security measures. Genetic algorithms can develop a wide range of countermeasure strategies, from access control policies to intrusion detection rules, providing comprehensive protection against cyber threats.
@@ III-C Strategy Objectives
  • Minimizing risk: This objective reflects the intention to reduce exposure to potential risks or adverse events. Such an objective indicates a significant concern to manage any uncertainties that may exist and to ensure the stability and reliability of the architecture. To achieve this goal, organizations must implement practical functions within a structured risk assessment process. This process begins with the identification and classification of all critical information assets. Potential risks are then identified, using tools such as threat intelligence platforms. Vulnerabilities are analyzed to gauge their exploitability and potential impact. Risk is quantified by estimating the probability of each threat occurring and multiplying it by its potential impact. This method allows a quantitative assessment of each threat, which helps to prioritize risks.
  • Threat detection: This objective represents the need to identify potential threats. It emphasizes the importance of proactive measures and early detection of possible vulnerabilities. It is the objective presented in the IDS. Threat detection is equivalent to having a vigilant security guard constantly monitoring the premises for any signs of suspicious activity. It involves the deployment of sophisticated tools and techniques that act as digital sentries scanning networks and systems for any anomalous behavior or potential security breaches. If it detects an unauthorized attempt to access sensitive data or unusual patterns indicative of a cyber-attack, it immediately triggers an alert.
  • Minimizing risk and cost: The combination of these objectives means a dual focus on cost-effectiveness and risk mitigation. It implies a strategy aimed at balancing economic considerations with security concerns. This approach may include conducting comprehensive risk assessments to identify potential vulnerabilities and threats, followed by prioritizing mitigation efforts based on their cost-effectiveness and impact reduction. Organizations, for example, could invest in robust cybersecurity plans that deliver a high return on investment by effectively mitigating common threats while maintaining manageable costs.
  • Countermeasure standard proposal: This objective indicates the intention to develop and propose standardized countermeasures. At the same time, it suggests a proactive role in the configuration of best practices and defensive strategies in the field. Standardized approaches are of particular importance, as they offer the advantage of unification. By implementing a standard, information can be shared more directly, easily, and quickly between different response teams. In this way, it not only facilitates communication between teams but also streamlines the process of incident coordination and response, improving the efficiency and effectiveness of security management.
@@ III-D Measurements and results
  • Positive rate: Often referred to as a true positive rate. Is a performance metric that measures the proportion of true positive predictions out of all actual positive instances  [78]. Equation 1 shows the mathematical expression of this metric, where True Positives (TP) represent instances that a model correctly recognizes or predicts as belonging to the attribute being tested. Conversely, False Negatives (FN) represent instances of the positive class that were incorrectly identified as negative.
    $TP\_rate=\frac{TP}{TP+FN}$ (1)
@@ III-E Use Case
  • Critical Infrastructure: This value is applicable in the field of critical infrastructures. Critical infrastructures refer to the fundamental systems, assets, facilities, and networks that are vital for the functioning of a society and its economy [79, 80]. Such a value implies the relevance and effectiveness of the proposed system in safeguarding and optimizing the essential components of the infrastructure.
  • Internet of Things (IoT): This value emphasizes the adaptability and functionality of a proposed system into the interconnected IoT ecosystem. IoT refers to a vast network of interconnected devices, objects, and systems that can communicate and exchange data with each other over the internet [81, 82].
@@ III-F Features Summary
  • Strategy Objectives where seven values are recognized: Minimizing risk, Vulnerability assessment, Minimizing risk and cost, Threat Detection, Calculation of cost-damage probabilities in attacks, Mitigate attack and Countermeasure standard proposal.
@@ IV Survey
For each of the papers a description of the research proposal is conducted. Moreover, an analysis and conclusions of the innovations are provided below highlighting the advantages and/or disadvantages of each of the approaches. Furthermore, this analysis is based on the features selected in Section III, ensuring that each paper will be discussed under the same conditions in terms of the proposed features.
@@ IV-A Cybersecurity risk analysis model using fault tree analysis and fuzzy decision theory [83]
In [83], the authors present a model that integrates Fault Tree Analysis (FTA) [84], decision theory [33], and fuzzy theory [85] to address two key objectives. FTA provides a structured and systematic way to identify and evaluate potential system failures by modeling the relationships between various components or events within a system. Decision theory is a field that encompasses various methods and tools for making rational and informed decisions. Fuzzy theory is a mathematical framework that allows for the representation of uncertainty and vagueness in a systematic manner.
Particularly, the primary objective of FTA is to identify the minimal cut set, which represents the minimum combination of basic events leading to the occurrence of the top event, for example the combination: “CyberAttack - Fault Information System -Unauthorized Access”. By analyzing these cut sets, appropriate actions can be prioritized and implemented to prevent the manifestation of the top event and identify vulnerabilities within the system. To comprehensively assess the complexity, magnitude, and impact of a cyber-attack, this research introduces the utilization of fuzzy theory and decision-making techniques under conditions of uncertainty.
@@
The third phase, focuses on addressing the uncertainties involved in risk analysis in the context of cybersecurity. The paper proposes the use of models presented in [86] and emphasizes the need to consider multiple criteria for evaluating the consequences of a cybersecurity solution. Specifically, the two criteria used in this paper are financial losses and time for restoration. In this sense, the importance of considering cost implications and the ability to repair or replace damaged services quickly is highlighted. The fourth phase focuses on the fuzzy assessment of potential accidents. During this stage, managerial experts evaluate alternative options using the identified criteria. Finally, the last phase involves aggregating all of the criteria used and ordering the alternatives, according to the magnitude of their consequences and using custom equations. These equations are designed for representing maximum risk using an estimated consequence of actions under nature states. The purpose of this representation is used to adopt countermeasures that minimize the risk identified.
@@
The initial focus of the research is on the data representation used for the countermeasure selection in the cybersecurity context. To achieve this objective, the authors define specific information objects that serve as input for the process. These objects are vulnerabilities, products, weaknesses, exploits, attack patterns, configurations, and countermeasures. The relationships between these objects are established through the analysis of open security databases such as Snyk [88] or Mend [89]. In this manner, events can be associated with these objects to enhance their representation and enable more accurate utilization of the information. The data collected by the system can be categorized into two types: input and output. Input data includes external information such as vulnerabilities, code weaknesses, attack patterns, and network analysis results, while output data comprises the results of attack modeling and analysis, including AGs, security metrics, impact assessment, chosen countermeasures, and identified weaknesses. The system utilizes both input and output data to enhance its understanding of the security landscape and enable effective decision-making.
After introducing the data representation, the research presents the proposed model for analyzing attacks. This model follows several steps: i) constructing an AG based on existing and potential vulnerabilities, ii) identifying malicious actions using logs and alerts, iii) generating an attack subgraph to depict the sequence of actions by the attacker, iv) integrating potential attack scenarios and countermeasures, v) calculating security metrics for risk analysis, and vi) presenting the resulting information for selecting appropriate responses.
@@ IV-C Heuristic Approach Towards Countermeasure Selection using Attack Graphs [91]
In [91], an approach of heuristic search is followed to select the optimal countermeasure deployment under a given budget constraint. The approach implements AGs to model the risk present in the system. In order to do this, the proposed system considers vulnerability exploitation consequences and examines potential attacker lateral movements. Additionally, it represents the network’s physical topology as well as the vulnerabilities of the network protocols to minimize the system risk within the budgetary limitations.
@@
The solution relies on the shortest path problem where the A* heuristic algorithm [75] is used to find the most optimal set of countermeasures efficiently. The novelty of the article relies upon using a particular risk equation, which indicates the presented risk of exploiting an asset provided by CVSS scores, together with the associated countermeasure, thus allowing the evaluation of the risk of the system without regenerating the AG, saving computational time. MulVAL [92] is used as a security analyze based on logic, serving as a tool for the analysis of AGs. This tool uses the Prolog [93] language to represent the pre and post-conditions for exploiting a vulnerability and then identify the necessary countermeasure. The authors define countermeasures as security products characterized by their implemented defense mechanisms, such as firewalls or antivirus software, and the associated mitigation actions, like disabling or enabling ports. The proposed system begins with a risk assessment using AGs. Relevant countermeasures are then identified by matching them with mitigation actions at AG nodes and determining their network execution points. Risk calculation follows using equations derived from AG preprocessing. The A* algorithm is employed to find the shortest path, minimizing risk while considering cost, and yielding an optimal countermeasure plan. Evaluation involves assigning costs for countermeasure deployment, with the algorithm adjusting the budget based on risk reduction. As the budget increases, system risk decreases, but algorithm execution time may rise due to consideration of additional, potentially costlier countermeasures.
@@
As mentioned before, this paper presents a methodology based on AGs for the purpose of risk minimization of a system. For this purpose, an A* algorithm is used for the selection of countermeasures. Analyzing the paper, a disadvantage identified concerns the limitations that conflict with the policies of the organization in which the system is implemented. In other words, the algorithm does not take into account certain limitations when selecting countermeasures. That is, in the risk graph shown in the paper, the higher the budget, the more the risk is reduced, even to 0%. This is however not useful in the real world, since the algorithm could propose countermeasures that disable relevant services or communications. Nevertheless, the use of MulVAL for AG security analysis is an interesting point. Furthermore, it alludes to establishing a maximum cost, on which the heuristic-based algorithm is built, which increases the realism of the design. Therefore, it is easy to spot that this is a very interesting proposal that, once the possibility of taking into account certain requirements is added, can be very relevant.
@@ IV-D Introducing Deep Learning Self-Adaptive Misuse Network Intrusion Detection Systems [94]
Due to the large number of interconnected devices that may exist in a network, the ability to handle their capabilities is typically beyond human limits. In the same way, IDSs are unable to automatically adapt to the many changes that occur in their environment, which reduces their efficiency. For example, by inserting a new device into the network it may generate services or communications or even vulnerabilities never seen in the network before. To solve the aforementioned challenges, in [94], Deep Learning (DL) is proposed to identify unknown threats in unknown environments. In particular, a self-taught learning methodology is used, an ML methodology that can train itself using unlabeled data to improve a supervised classification problem. This methodology is integrated within the MAPE-K structure [95], which is based on the steps Monitor (M), Analyze (A), Plan (P), and Execute(E) while using a Knowledge (K) base as support. The proposed system is based on environmental states and network flows. Network changes are seen as potential incidents that may contain attacks or vulnerabilities. For this purpose, network flows are used, which, by possessing relevant information, can identify unknown and unseen network flows that may indicate the presence of previously unknown attacks.
In the monitor phase, the sensors of the system are coordinated to acquire the basic knowledge of the network, as well as network mappers [96] generate a network inventory. Identifying the network characteristics alteration events that require activating the IDS. In parallel, network sniffers are used to capture network flows through the detection engine of the IDS. In this phase, network traffic is stored in the Knowledge base for later phases. In the Analyze phase, an analyzer leverages network audit tools to generate structured network flows from network traffic that can bring various features. These features are fed into the IDS engine, which will in turn use a supervised model for detection purposes and, on the other hand, it will be used as unlabeled data for the self-taught learning adaptive process.
Up to this point, the Monitor and the Analyzer identified environmental changes in the network, while the Knowledge component consolidated the network flows generated at the time the changes occurred. At this moment, the IDS may face unknown network traffic, which is why the planning phase undertakes to utilize unsupervised feature learning techniques. This technique is an NN that applies back-propagation called sparse auto-encoder [97] to learn sparse representations of the input data. The output of the planning phase is a feedforward auto-encoder that is used to reconstruct the initial set of labeled data into a new training dataset for a supervised learning algorithm. After training the new model, the old model, which due to environmental changes had started to have efficiency problems, can be replaced.